what equation could help me solve this problem??
Briana and her sister Molly both want to buy the same model bicycle. Briana needs $73 more before she can afford the bike. Molly needs $65 more. If they combine their money, they will have just enough to buy one bicycle that they could share. What is the cost of the bicycle?

Answer by ankor@dixie-net.com(22740)

what equation could help me solve this problem??
Briana and her sister Molly both want to buy the same model bicycle. Briana needs $73 more before she can afford the bike. Molly needs $65 more. If they combine their money, they will have just enough to buy one bicycle that they could share. What is the cost of the bicycle?
:
Let x = cost of bike
:
It says,"Briana needs $73 more before she can afford the bike.", so we can say:
(x - 73) = B, $amt
:
" Molly needs $65 more."
(x - 65) = M's $amt
:
The equation: B's amt + M's amt = cost of bike
(x - 73) + (x - 65) = x
2x - 138 = x
2x - x = 138
x = $138